OpenStudy (anonymous):

The perimeter of a rectangle is 28 ft. the length is 6 ft longer than the width, find the dimensions? The formula for the perimeter, P, of a rectangle is 2L + 2W = P Since the length is 6 feet longer, W = L - 6.

OpenStudy (anonymous):

l=6 ft longer than w so l=6+w 28=2w+2(6+w) 28=2w+12+2w 28=4w+12 16=4w 4=w l=6+w l=10 28=2(10)+2(4) 28=20+8 28=28

OpenStudy (anonymous):

So if I were to write the first equation would it be 2L+2W= 28 then the second equation l=w+6?? would it be 28=4L-12 +12 +12 40=4L L= 10 ?